In a resonance tube experiment, a vibrating fork of frequency 256 Hz gave resonance when the water level was 30.0 cm below the open end of the tube. If the next position of resonance was 100 cm, what is the velocity of sound in air?[
1
Expert's answer
2020-05-25T11:18:13-0400

The velocity of sound in air is


"v=2f(L_2-L_1)=2\cdot256\cdot(1-0.3)=358\text{ m/s}."